A scale model of a barn is 1.2m tall and 0.8 wide. The width of the actual barn is 20 m . What is the actual height of the barn?
Answer by macston(5194)

x=actual height
20m/0.8m=x/1.2m Multiply each side by 1.2m
1.2m(20m)/0.8m=x
24sq m/0.8m=x
30m=x ANSWER: The actual height of the barn is 30 meters.